OpenStudy (anonymous):

solve with substitution y=3x-1 y=-5x+3

OpenStudy (anonymous):

You can just set the two equations equal to each other, since you have y. 3x-1 = -5x +3 so x = .5 Then y = 3x - 1 y =3*.5 -1 y = 1.5-1 y = .5 So x =.5, y =.5